OpenStudy (anonymous):

Consider the function f(x) = 2sin(x2) on the interval 0 ≤ x ≤ 3. (a) Find the exact value in the given interval where an antiderivative, F, reaches its maximum. x = (b) If F(1) = 9, estimate the maximum value attained by F. (Round your answer to three decimal places.) y ≈

OpenStudy (inkyvoyd):

set 2sin(x^2)=0

OpenStudy (inkyvoyd):

Wait, over an interval -_-

OpenStudy (inkyvoyd):

Ignore the interval, and just solve that trig equation.

OpenStudy (anonymous):

i know that the answer for part "a" is

OpenStudy (anonymous):

square(pi)

OpenStudy (inkyvoyd):

square or sqrt?

OpenStudy (anonymous):

i dont know how to do part b

OpenStudy (anonymous):

sqrt

OpenStudy (inkyvoyd):

Well, your interval s defined only up to 3. Is that supposed to be that way, or is the three suposed to be pi?

OpenStudy (anonymous):

up to 3

OpenStudy (anonymous):

a)\[2*\sin(2*x)=0\rightarrow x=k*pi/2\] In this interval x=pi/2

OpenStudy (inkyvoyd):

Do you understand a though?

OpenStudy (anonymous):

well i got x= 

OpenStudy (inkyvoyd):